OpenStudy (anonymous):

Ab is tangent to circle O at A. IF AO=30 and BC=48 what is AB?

OpenStudy (anonymous):

Please see your question again .

OpenStudy (anonymous):



OpenStudy (anonymous):

72?

OpenStudy (anonymous):

\[=\sqrt{\left( 78+30 \right)\left( 78-30 \right)}=\sqrt{108*48}=\sqrt{9*4*3*16*3}=3*2*4*3=72\] You are correct.

OpenStudy (anonymous):

\[\frac{ x }{ 12 }=\frac{ 21 }{ 9 },x=\frac{ 21 }{ 9 }\times12=?\]

OpenStudy (anonymous):

28

OpenStudy (anonymous):

correct